Heidi makes a bank deposit of $59500 with 96 bills consisting of $500 and $1000 bills. How many of each kind of bills did she deposit? , Im very Thankful that u had this site.. Tnx a lot!!!

Answer by ankor@dixie-net.com(22740)

Heidi makes a bank deposit of $59500 with 96 bills consisting of $500 and $1000 bills.
How many of each kind of bills did she deposit?
:
Let t = no. of $1000 bills
Let f = no. of $500 bills
:
A total of 96 bills, therefore
t + f = 96
:
A value of $59500
1000t + 500f = 59500
:
We can use elimination here. multiply the 1st equation by -500, add to the above equation
1000t + 500f = 59500
-500t - 500f =-48000
-------------------------additions eliminates f, find t
500t = 11500
t = 11500/500
t = 23 ea $1000 bills
:
Find f using the 1st equation
23 + f = 96
f = 96-23
f = 73 ea $500 bills
:
:
You can check these solutions using the value equation: